# Experimental aspects of the Quantum Tomography of tau lepton pairs at a Higgs factory collider

### 13 Mar 2026 v1

@@ Abstract
Quantum Tomography of tau lepton pairs produced at a Higgs Factory collider will enable measurements of their spin correlations arising from quantum entanglement. Such measurements rely on the ability to measure the components of and correlations between the taus’ spins. We present a method to fully reconstruct the kinematics of tau pair events at an electron-positron Higgs factory collider, making use of measured particles’ momenta and impact parameters. This procedure results in several consistent solutions per event, which can be assigned weights according to various event properties. Full kinematic reconstruction allows the optimal extraction of the taus’ spin orientation via polarimeters. We estimate the precision with which polarimeters can be reconstructed in an ideal detector, and quantify the effects of more realistic detector performance. We conclude that for this analysis, achieving a photon angular resolution of around 1 mrad is the most crucial aspect of detector performance, while photon energy resolution and vertex detector performance are significantly less important.
@@ 2 Kinematic reconstruction
We apply this method to $e^{+}e^{-}\to\tau^{+}\tau^{-}(\gamma)$ events at the ILC at a CoM energy of 250 $\mathrm{GeV}$. Events were simulated using the WHIZARD event generator [20], modeling the effects of both ISR and the expected colliding beam energy spectrum at ILC, smeared due to both the beam energy spread and beamstrahlung. Taus were decayed by TAUOLA [21]. The analysed samples correspond to $40~\mathrm{pb^{-1}}$ of unpolarised collisions, close to 0.4 M events in total. We note that in these samples care was taken to accurately model correlations between the longitudinal tau spin components, however correlations between other components may less reliable. This does not affect the results of this paper, which concern the precision with which these components can be extracted. The simulated distribution of the $\tau^{-}\tau^{+}$ invariant mass is shown in Fig. 2. The population of events is enhanced around the nominal CM energy (250 $\mathrm{GeV}$), the Z boson mass, and around twice the tau mass (dominated by photon exchange).
@@ 4 Polarimeter extraction
$\mathbf{h}(\tau^{\pm}\rightarrow\pi^{\pm}\pi^{0}\nu)\propto$ $m_{\tau}(E_{\pi^{\pm}}-E_{\pi^{0}})(\mathbf{p}_{\pi^{\pm}}-\mathbf{p}_{\pi^{0}})$ (13)
$+\frac{1}{2}(p_{\pi^{\pm}}+p_{\pi^{0}})^{2}\mathbf{p}_{\nu},$
where $p_{\pi^{\pm}}$, $p_{\pi^{0}}$, $p_{\nu}$ are respectively the four-momenta of the charged and neutral pions, and of the neutrino, defined in the tau rest frame [22]. For the more complex three-pion decays $\tau^{\pm}\to\pi^{\pm}\pi^{0}\pi^{0}\nu$ and $\tau^{\pm}\to\pi^{\pm}\pi^{\pm}\pi^{\mp}\nu$ we use the method and associated computer code of [23].

### 28 Apr 2026 v2

@@ Abstract
Quantum Tomography of tau lepton pairs produced at a Higgs Factory collider will enable measurements of their spin correlations arising from quantum entanglement. Such measurements rely on the ability to measure the components of and correlations between the taus’ spins. We present a method to fully reconstruct the kinematics of tau pair events at an electron-positron Higgs factory collider, making use of measured particles’ momenta and impact parameters. This procedure results in several consistent solutions per event, which can be assigned weights according to various event properties. Full kinematic reconstruction allows the optimal extraction of the taus’ spin orientation via polarimeters. We estimate the precision with which polarimeters can be reconstructed in an ideal detector, and quantify the effects of more realistic detector performance. We conclude that for this analysis, achieving a photon angular resolution of around 0.1 mrad is the most crucial aspect of detector performance, while photon energy resolution and vertex detector performance are significantly less important.
@@ 2 Kinematic reconstruction
We apply this method to $e^{+}e^{-}\to\tau^{+}\tau^{-}(\gamma)$ events at the ILC at a CoM energy of 250 $\mathrm{GeV}$. Events were simulated using the WHIZARD event generator [20], modeling the effects of both ISR and the expected colliding beam energy spectrum at ILC, smeared due to both the beam energy spread and beamstrahlung. Taus were decayed by TAUOLA [21, 22]. The analysed samples correspond to $40~\mathrm{pb^{-1}}$ of unpolarised collisions, close to 0.4 M events in total. We note that in these samples care was taken to accurately model correlations between the longitudinal tau spin components, however correlations between other components may less reliable. This does not affect the results of this paper, which concern the precision with which these components can be extracted. The simulated distribution of the $\tau^{-}\tau^{+}$ invariant mass is shown in Fig. 2. The population of events is enhanced around the nominal CM energy (250 $\mathrm{GeV}$), the Z boson mass, and around twice the tau mass (dominated by photon exchange).
@@ 4 Polarimeter extraction
$\mathbf{h}(\tau^{\pm}\rightarrow\pi^{\pm}\pi^{0}\nu)\propto$ $2(q\cdot p_{\nu})\mathbf{q}-q^{2}\mathbf{p}_{\nu},$ (13)
where $p_{\pi^{\pm}}$, $p_{\pi^{0}}$, $p_{\nu}$ are respectively the four-momenta of the charged and neutral pions, and of the neutrino defined in the tau rest frame, and $q=p_{\pi^{\pm}}-p_{\pi^{0}}$ [21]. For the more complex three-pion decays $\tau^{\pm}\to\pi^{\pm}\pi^{0}\pi^{0}\nu$ and $\tau^{\pm}\to\pi^{\pm}\pi^{\pm}\pi^{\mp}\nu$ we use the method and associated computer code of [23].
@@ 6 Conclusions and outlook
Acknowledgements
I thank my colleagues in the ILD software group for generating the MC samples used in this study, M. Berggren for assistance with running SGV, and A. Korchin for pointing out an error in the form of the $\tau^{\pm}\rightarrow\pi^{\pm}\pi^{0}\nu$ polarimeter (Eq. 13) given in an earlier version of this paper.